According to the collision theory model,why does temperature affect the value of the rate constant.
1
Expert's answer
2014-01-25T09:40:55-0500
According to collision theory, the reaction rate constant can be expressed as follows:
k = z0*exp(-Ea/RT)
where z0 is the collision frequency, Ea - activation energy of the process.
z0 = N*sigma*is the reaction cross section, mu is the reduced mass of two reacting particles:
1/mu = 1/m1 + 1/m2
The temperature is included in this equation twice: firstly it exists under exponential; secondly it is included into the collision frequency z0 in power 1/2.